### v2

@@ 4 Model 2: many-body probe (half filling)
where $\Delta$ represents the strength of the interaction between particles and $\sigma_{l}^{x,y,z}$ are the Pauli matrices acting on site $l$. In the single excitation subspace and with $\Delta{=}0$, the Hamiltonian in Eq. (4) reduces to the single-particle case in Eq. (4). We consider the time evolution of a system initially prepared in a Neel state, i.e. $\mathinner{|{\Psi(0)}\rangle}=\mathinner{|{1,0,\cdots,1,0}\rangle}$ for even $L$ and $\mathinner{|{\Psi(0)}\rangle}=\mathinner{|{1,0,\cdots,0,1}\rangle}$ for odd $L$. More details about the time evolution and complex pattern of many-body Bloch oscillations are available in Appendix B. We first put our focus on $\Delta{=}0$ (non-interacting case) and $\Delta{=}1$ (isotropic Heisenberg Hamiltonian). The results for the more general case of $10{<}\Delta{<}$ 1 are presented and discussed in the next section. In Figs. 3(a) and 3(b), we plot the long time limit of $\mathcal{F}_{Q}/t^{2}$ as a function of $h/J$ for various system sizes $L$ for $\Delta{=}0$ and $\Delta{=}1$, respectively. In both interacting and non-interacting regimes, the normalized QFI peaks at the transition point $h_{c}$ and it shows algebraic decay of the form $\mathcal{F}_{Q}/t^{2}\sim h^{-2}$ in the localized phase, which is similar to the single excitation case. However, unlike the single excitation scenario, even in the localized phase the QFI shows size dependence. Furthermore, one can see that in the presence of interaction, the behavior of the QFI qualitatively changes as: (i) it shows fluctuations in the vicinity of the transition point; and (ii) its value becomes smaller than the non-interacting case (i.e. $\Delta=0$). To extract the scaling with system size, in Figs. 3(c) and 3(d) we plot the long-time value of $\mathcal{F}_{Q}/t^{2}$ at $h{=}h_{c}$ as a function of system size $L$ for $\Delta{=}0$ and $\Delta{=}1$, respectively. Interestingly, in both cases, the scaling suggests $\beta>2$, which indicates a super-Heisenberg scaling. Especially, in the non-interacting case $\Delta=0$, it takes the form of $\mathcal{F}_{Q}(h_{c})\sim t^{2}L^{3}$. The difference of $\beta>2$ here, compared to the single-particle result of $\beta{=}2$ at the transition point, is related to the number of excitations that will be further explained in the next section.
@@ 5 Effect of excitations
In the previous section, we focused on the half-filling case, namely $N{=}L/2$. In this section, we investigate the impact of the number of excitations in more details. For the sake of symmetry, we consider odd system sizes $L$ and choose the initial states to be symmetric around the center of the chain. For instance, if the system size is $L{=}7$ then the initial state with $N=2$ and $N=3$ excitations are set to $\mathinner{|{0,0,1,0,1,0,0}\rangle}$ and $\mathinner{|{0,1,0,1,0,1,0}\rangle}$, respectively. In Figs. 4(a) and 4(b) the normalized QFI is plotted for different excitation numbers $N$ in a system of size $L{=}17$ for $\Delta=0$ and $\Delta=1$, respectively. Notably, increasing the excitations always enhances the QFI, indicating better sensitivity. In addition, in the localized phase the QFI shows excitation-dependence. In order to discriminate the QFI enhancement due to the system size from the excitation number, in Figs. 4(c) and 4(d) we plot the long-time behavior of $\mathcal{F}_{Q}/t^{2}$ as a function of $h/J$ for a fixed number of excitation $N{=}3$ for various choices of $L$. Interestingly, the size dependence completely disappears for $h>h_{c}$ indicating that the sensitivity is only enhanced through the number of excitations when the system is in the localized phase (The scaling of the normalized QFI at the transition points in Fig. 4(c) and (d) are discussed in Appendix C). These observations suggest the following ansatzes for the QFI in many-body probes with multiple excitations
@@
where $\alpha$ is the exponent which quantifies the dependence on the excitation number. In Fig. 5(a) for a fixed system of size $L{=}17$, we plot the long-time value of $\mathcal{F}_{Q}/t^{2}$ as a function of $N$ which clearly shows an algebraic growth. To complete the analysis, in Fig. 5(b) the exponent $\alpha$ is depicted as a function of $\Delta$, for two choices of $h$, one in the localized phase, i.e. $h/J{=}5$, and one at the transition point $h{=}h_{c}$. As the figure shows, while increasing the excitations always enhances the QFI, its rate of enhancement decreases by increasing the interaction strength $\Delta$. This is an interesting observation which can be related to various configurations that the system gets due to the interaction term $\Delta$ and effectively enhances the localization [103]. Finally, it is worth emphasizing that in the case of half-filling, namely $N{=}L/2$, and in the range of $h\leq h_{c}$ (i.e. extended phase) Eq. (7) suggests that the QFI scales as $\mathcal{F}_{Q}\sim t^{2}L^{\beta+\alpha}$. The main advantage of Eq. (7) is to separate the dependence on $L$ and $N$. Therefore, one can extract $\beta$ from the single excitation analysis, shown in Fig. 2(c), and the exponent $\alpha$ can be read from Fig. 5(b). For instance, at the transition point $h{=}h_{c}$ the single excitation analysis gives $\beta{=}2$ and from Fig. 5(b) one reads $\alpha{=}0.88$ at $\Delta{=}0$ and $\alpha=0.63$ at $\Delta{=}1$. These lead to $\mathcal{F}_{Q}\sim t^{2}L^{2.88}$ and $\mathcal{F}_{Q}\sim t^{2}L^{2.63}$ for $\Delta{=}0$ and $\Delta{=}1$, respectively. These scaling functions are very close to the fittings directly extracted in Fig. 3.
@@ 6 Practical implementation
As discussed before, QFI quantifies the ultimate precision limit in a sensing procedure which can only be achieved by an optimal measurement. In practice, however, due to practical constraints performing such an optimal measurement might be very challenging. In this section, we consider simple available measurements in both single- and multi-particle cases. For the single excitation sensors, we consider a simple position measurement described by projectors $\{\pi_{l}{=}\mathinner{|{l}\rangle}\mathinner{\langle{l}|}\}$, where $l{=}1,2,\cdots,L$. Therefore, the probability of each outcome is $p_{l}(t){=}|\braket{l}{\Psi(t)}|^{2}$. The obtainable precision by this measurement is bounded by the CFI, see Eq. (1). In Fig. 6(a), we compare the long-time behavior of the CFI and the QFI in a single-excitation probe of size $L=16$ as a function of $h/J$. As the figure shows, the CFI qualitatively follows the QFI, though it always takes lower values. In the inset, we plot the ratio $\mathcal{F}_{C}/\mathcal{F}_{Q}$ as a function of $h/J$ which shows that our simple measurement is almost optimal in the extended phase (i.e. $\mathcal{F}_{C}/\mathcal{F}_{Q}\simeq 1$) and it becomes sub-optimal (i.e. $\mathcal{F}_{C}/\mathcal{F}_{Q}\sim 0.5$) as localization strength is enhanced. Interestingly, the lowest ratio of $\mathcal{F}_{C}/\mathcal{F}_{Q}$ takes place around the transition point showing that our simple measurement scheme cannot efficiently extract information from the complex critical quantum state.
In the case of a multi-particle probe, we measure the configuration of particles which is obtained by a simple particle detection imaging. The measurement can be described by $\{\pi_{{}_{\mathbf{Z}}}{=}\mathinner{|{\mathbf{z}}\rangle}\mathinner{\langle{\mathbf{z}}|}\}$, where $\mathinner{|{\mathbf{z}}\rangle}$ accounts for all possible particle configurations in the probe. For instance, in a probe of size $L$ with $N$ particles (i.e. excitations), there are (${L\choose N}$ different configurations. In the case of $N{=}1$, the problem reduces to the position measurement, discussed above. Each configuration is observed by the probability $p_{{}_{\mathbf{z}}}(t){=}|\braket{\mathbf{z}}{\Psi(t)}|^{2}$, for which one can obtain the corresponding CFI. In Fig. 6(b), we compare the long-time behavior of the CFI and the QFI in a half-filled probe of size $L=16$ (with $N=L/2$ excitations) initialized in the Neel state as a function of $h/J$. Similar to the case of single excitation, the CFI qualitatively follows the QFI, though with lower values. In the inset, we plot $\mathcal{F}_{C}/\mathcal{F}_{Q}$ as a function of $h/J$. Similar to the single excitation case, it shows that a simple configuration measurement is almost optimal in the extended phase and becomes sub-optimal with $\mathcal{F}_{C}/\mathcal{F}_{Q}\simeq 0.5$ as $h/J$ increases. Again the lowest ratio of $\mathcal{F}_{C}/\mathcal{F}_{Q}$ is obtained around the transition point where our simple measurement scheme cannot fully extract the information from the complex critical quantum state of the system.

### v3

@@ 4 Model 2: many-body probe (half filling)
where $\Delta$ represents the strength of the interaction between particles and $\sigma_{l}^{x,y,z}$ are the Pauli matrices acting on site $l$. In the single excitation subspace and with $\Delta{=}0$, the Hamiltonian in Eq. (4) reduces to the single-particle case in Eq. (4). We consider the time evolution of a system initially prepared in a Neel state, i.e. $\mathinner{|{\Psi(0)}\rangle}=\mathinner{|{1,0,\cdots,1,0}\rangle}$ for even $L$ and $\mathinner{|{\Psi(0)}\rangle}=\mathinner{|{1,0,\cdots,0,1}\rangle}$ for odd $L$. More details about the time evolution and complex pattern of many-body Bloch oscillations are available in Appendix B. We first put our focus on $\Delta{=}0$ (non-interacting case) and $\Delta{=}1$ (isotropic Heisenberg Hamiltonian). The results for the more general case of $10{<}\Delta{<}$ 1 are presented and discussed in the next section. In Figs. 3(a) and 3(b), we plot the long time limit of $\mathcal{F}_{Q}/t^{2}$ as a function of $h/J$ for various system sizes $L$ for $\Delta{=}0$ and $\Delta{=}1$, respectively. In both interacting and non-interacting regimes, the normalized QFI peaks at the transition point $h_{c}$ and it shows algebraic decay of the form $\mathcal{F}_{Q}/t^{2}\sim h^{-2}$ in the localized phase, which is similar to the single excitation case. However, unlike the single excitation scenario, even in the localized phase the QFI shows size dependence. Furthermore, one can see that in the presence of interaction, the behavior of the QFI qualitatively changes as: (i) it shows fluctuations in the vicinity of the transition point; and (ii) its value becomes smaller than the non-interacting case (i.e. $\Delta=0$). To extract the scaling with system size, in Figs. 3(c) and 3(d) we plot the long-time value of $\mathcal{F}_{Q}/t^{2}$ at $h{=}h_{c}$ as a function of system size $L$ for $\Delta{=}0$ and $\Delta{=}1$, respectively. In both cases, the scaling suggests $\beta>2$, which indicates a quantum-enhanced scaling. Especially, in the non-interacting case $\Delta=0$, it takes the form of $\mathcal{F}_{Q}(h_{c})\sim t^{2}L^{3}$. The difference of $\beta>2$ here, compared to the single-particle result of $\beta{=}2$ at the transition point, is related to the number of excitations that will be further explained in the next section.
(a)Refer to caption(b)
Refer to caption(c)Refer to caption(d)
It is worth emphasizing that the transition point $h_{c}$ decreases as the system size increases, see Fig. 2(a) and Figs. 3(a)-(b). In other words, in the thermodynamic limit the ergodic regime over which the quantum-enhanced sensitivity can be obtained shrinks to a vanishingly small region. This has an implication for the scaling of the QFI. For any given $h$, as the system size increases, at some point the system enters the localized phase and hence the QFI becomes size independent. For instance, in the case of the single-particle probes, for any given $h$ when the system size exceeds $L_{c}{\simeq}8J/h$ the probe becomes localized and the QFI saturates. This means that quantum-enhanced sensitivity can only be observed over finite sizes, which is in agreement with the results of Ref. [106] that at best predicts Heisenberg scaling (i.e. $\mathcal{F}_{Q}\sim t^{2}L^{2}$) for single-body interacting systems, such as our Stark Hamiltonian. Note that although the quantum-enhanced sensitivity does not survive in the thermodynamic limit, in practice, all quantum probes are essentially finite and thus their scaling over short system sizes matters more than the true thermodynamic limit.
@@ 5 Effect of excitations
In the previous section, we focused on the half-filling case, namely $N{=}L/2$. In this section, we investigate the impact of the number of excitations in more details. For the sake of symmetry, we consider odd system sizes $L$ and choose the initial states to be symmetric around the center of the chain. For instance, if the system size is $L{=}7$ then the initial state with $N=2$ and $N=3$ excitations are set to $\mathinner{|{0,0,1,0,1,0,0}\rangle}$ and $\mathinner{|{0,1,0,1,0,1,0}\rangle}$, respectively. In Figs. 4(a) and 4(b) the normalized QFI is plotted for different excitation numbers $N$ in a system of size $L{=}17$ for $\Delta{=}0$ and $\Delta{=}1$, respectively. Notably, increasing the excitations always enhances the QFI, indicating better sensitivity. In addition, in the localized phase the QFI shows excitation-dependence. In order to discriminate the QFI enhancement due to the system size from the excitation number, in Figs. 4(c) and 4(d) we plot the long-time behavior of $\mathcal{F}_{Q}/t^{2}$ as a function of $h/J$ for a fixed number of excitation $N{=}3$ for various choices of $L$. The scaling of the normalized QFI at the transition points in Fig. 4(c) and (d) are discussed in Appendix C. These observations suggest the following ansatzes for the QFI in many-body probes with multiple excitations
@@
where $\alpha$ is the exponent which quantifies the dependence on the excitation number. The size dependence completely disappears for $h>h_{c}$ when the system enters the localized phase. This can be understood as every excitation in the system can only fluctuates locally and thus dependence on the system size disappears. In this situation, the sensitivity is only enhanced by either waiting for longer times $t$ or increasing the number of excitations $N$. The reason behind this is that every excitation localizes at some spatial part of the probe where it fluctuates locally. Every small changes from each of these excitations contributes in enhancing the QFI of the system.
@@
In Fig. 5(a) for a fixed system of size $L{=}17$, we plot the long-time value of $\mathcal{F}_{Q}/t^{2}$ as a function of $N$ which clearly shows an algebraic growth. To complete the analysis, in Fig. 5(b) the exponent $\alpha$ is depicted as a function of $\Delta$, for two choices of $h$, one in the localized phase, i.e. $h/J{=}5$, and one at the transition point $h{=}h_{c}$. As the figure shows, while increasing the excitations always enhances the QFI, its rate of enhancement decreases by increasing the interaction strength $\Delta$. This is an interesting observation which can be related to various configurations that the system gets due to the interaction term $\Delta$ and effectively enhances the localization [107]. Finally, it is worth emphasizing that in the case of half-filling, namely $N{=}L/2$, and in the range of $h\leq h_{c}$ (i.e. extended phase) Eq. (7) suggests that the QFI scales as $\mathcal{F}_{Q}\sim t^{2}L^{\beta+\alpha}$. The main advantage of Eq. (7) is to separate the dependence on $L$ and $N$. Therefore, one can extract $\beta$ from the single excitation analysis, shown in Fig. 2(c), and the exponent $\alpha$ can be read from Fig. 5(b). For instance, at the transition point $h{=}h_{c}$ the single excitation analysis gives $\beta{=}2$ and from Fig. 5(b) one reads $\alpha{=}0.88$ at $\Delta{=}0$ and $\alpha=0.63$ at $\Delta{=}1$. These lead to $\mathcal{F}_{Q}\sim t^{2}L^{2.88}$ and $\mathcal{F}_{Q}\sim t^{2}L^{2.63}$ for $\Delta{=}0$ and $\Delta{=}1$, respectively. These scaling functions are very close to the fittings directly extracted in Fig. 3.
6 Source of the quantum-enhanced sensitivity in the Stark probe
While the above analysis relies solely on numerical simulation, one can also shed light on the nature of quantum-enhanced sensitivity by focusing on the analytical bounds. In particular, one can decompose the Hamiltonians of Eq. (4), for a single excitation probe, and Eq. (4), for the many-body probe, into a general form of $H=H_{1}+hH_{2}$, where $H_{1}$ accounts for the hopping term and $H_{2}$ describes the gradient field. For a given initial state $\mathinner{|{\Psi(0)}\rangle}$ which is independent of $h$, the time evolution is given by $\mathinner{|{\Psi_{h}(t)}\rangle}=e^{-iHt}\mathinner{|{\psi(0)}\rangle}$. In Ref. [106], it has been shown that the QFI of the quantum state $\mathinner{|{\Psi_{h}(t)}\rangle}$ is fundamentally bounded by $\mathcal{F}_{Q}{\leq}t^{2}||H_{2}||^{2}$, where $||H_{2}||{=}\lambda_{\rm max}{-}\lambda_{\rm min}$ is the semi-norm with $\lambda_{\rm max}$ ($\lambda_{\rm min}$) being the maximum (minimum) eigenvalue of $H_{2}$. Although, the term $t^{2}||H_{2}||^{2}$ is only a bound for $\mathcal{F}_{Q}$, it provides several insights for the scaling of the QFI. First, it clearly demonstrates the quadratic scaling with respect to time. Second, it shows that the scaling with respect to system size $L$ is connected to the semi-norm of $H_{2}$.
In the case of a single excitation, see Eq. (4), the semi-norm $||H_{2}||{=}(L-1)$, resulting in the bound of $\mathcal{F}_{Q}{\leq}t^{2}(L-1)^{2}$. This fundamental bound shows that the QFI scaling exponent can only be $\beta{\leq}2$. Note that this is not inconsistent with Fig. 2(c) in which the extended phase supports $\beta{>}2$. In fact, as mentioned before, this region is only exist in finite system sizes as in the thermodynamic limit, where the critical point $h_{c}{\rightarrow}0$, the region shrinks to an infinitesimal region.
In the case of many-body probes with $N$ excitations, the semi-norm becomes $||H_{2}||{=}NL{-}N^{2}$. For the case of half filling where $N{=}L/2$, one gets $||H_{2}||{=}L^{2}{/}4$. This implies that $\mathcal{F}_{Q}{\leq}t^{2}L^{4}/16$. Indeed, our analysis in section 4 shows that $\mathcal{F}_{Q}{\sim}t^{2}L^{\beta+\alpha}$, in which $\beta{\leq}2$ comes from the single excitation analysis and $\alpha{\leq}1$ is extracted from Fig. 5(b). This means that the QFI scales as $\mathcal{F}_{Q}{\sim}t^{2}L^{3}$, which is clearly lower than the prediction of the analytic bound given by $t^{2}||H_{2}||^{2}{\sim}t^{2}L^{4}$.
7 Impact of decoherence
In practice, quantum systems are not isolated, and thus their dynamics cannot be described by unitary evolution. In fact, interaction with the environment can significantly affect the quality of quantum probes [108, 109]. To investigate this, we consider our quantum probe under the effect of dephasing. The dynamics is given by the Lindblad master equation
$\dot{\rho}=-i[H,\rho]+\gamma\sum_{l=1}^{L}\left(\mathcal{L}_{l}\rho\mathcal{L}^{\dagger}_{l}-\frac{1}{2}\{\mathcal{L}^{\dagger}_{l}\mathcal{L}_{l},\rho\}\right)\leavevmode\nobreak\ ,$ (9)
where $\rho$ is the density matrix of the probe, $\gamma$ is the dephasing rate, and $\mathcal{L}_{l}$ represents the Lindblad operators of site $l$. For pure dephasing, in general, the Lindblad operators are given by $\mathcal{L}_{l}{=}\sigma^{z}_{l}$. In the single-excitation probe, the Lindblad operators can also be written as $\mathcal{L}_{l}{=}\mathbb{I}-2\mathinner{|{l}\rangle}\mathinner{\langle{l}|}$, in which $\mathbb{I}{=}\sum_{l=1}^{L}\mathinner{|{l}\rangle}\mathinner{\langle{l}|}$. In Fig. 6(a), we plot the normalized QFI, namely $\mathcal{F}_{Q}/t^{2}$, at $h/J{=}0.1$ as a function of time for different values of dephasing rates $\gamma$ for the single-particle probe of size $L=16$. As the figure shows, by increasing $\gamma$ the QFI ultimately decays to zero and cannot grow quadratically in time. Similarly, for the multi-particle probe of size $L{=}6$ with $N{=}3$ excitations (i.e. half filling), we plot the normalized QFI for $h/J{=}0.5$ as a function of time for various choices of $\gamma$ in Fig. 6(b). The results show similar behavior with the single-excitation probe as the normalized QFI eventually goes to zero as time increases. This is because the steady state of the probe under dephasing dynamics is the featureless maximally mixed state which carries no information about the parameters of the system. One can also fix the dephasing rate and investigate the evolution of the QFI for different choices of $h$. In Figs. 6(c)-(d), we present the normalized QFI as a function of time for $\gamma/J=0.005$ and various choices of $h/J$ for the single- and multi-particle probes, respectively. Interestingly, by increasing $h$ the probe gets more localized and thus the action of dephasing becomes less destructive as the wave function is made of fewer components in its superposition.
(a)Refer to caption(b)
Refer to caption(c)Refer to caption(d)
Figure 6: The time evolution of $\mathcal{F}_{Q}/t^{2}$ for different values of $\gamma$ is plotted for: (a) the single-excitation probe with the system size $L=16$ at $h/J{=}0.1$; and (b) the half-filled excitation probe with the system size $L{=}6$ for $\Delta=0$ at $h/J{=}0.5$. The time evolution of $\mathcal{F}_{Q}/t^{2}$ for dephasing strength $\gamma/J{=}0.005$ and different values of $h/J$ is shown as a function of time for: (c) the single-excitation probe with the system size $L{=}16$; and (d) the half-filled excitation probe with the system size $L{=}6$ with $\Delta=0$.
@@ 8 Practical implementation
As discussed before, QFI determines a lower bound on the precision of an unbiased estimator given optimal measurements. In practice, however, performing such an optimal measurement might be very challenging. In this section, we consider simple available measurements in both single- and multi-particle cases. For the single excitation sensors, we consider a simple position measurement described by projectors $\{\pi_{l}{=}\mathinner{|{l}\rangle}\mathinner{\langle{l}|}\}$, where $l{=}1,2,\cdots,L$. Therefore, the probability of each outcome is $p_{l}(t){=}|\braket{l}{\Psi(t)}|^{2}$. The obtainable precision by this measurement is bounded by the CFI, see Eq. (1). In Fig. 7(a), we compare the long-time behavior of the CFI and the QFI in a single-excitation probe of size $L{=}16$ as a function of $h/J$. As the figure shows, the CFI qualitatively follows the QFI, though it always takes lower values. In the inset, we plot the ratio $\mathcal{F}_{C}/\mathcal{F}_{Q}$ as a function of $h/J$ which shows that our simple measurement is almost optimal in the extended phase (i.e. $\mathcal{F}_{C}/\mathcal{F}_{Q}\simeq 1$) and it becomes sub-optimal (i.e. $\mathcal{F}_{C}/\mathcal{F}_{Q}\sim 0.5$) as localization strength is enhanced. Interestingly, the lowest ratio of $\mathcal{F}_{C}/\mathcal{F}_{Q}$ takes place around the transition point showing that our simple measurement scheme cannot efficiently extract information from the complex critical quantum state.
@@
In the case of a multi-particle probe, we measure the configuration of particles which is obtained by a simple particle detection imaging. The measurement can be described by $\{\pi_{{}_{\mathbf{Z}}}{=}\mathinner{|{\mathbf{z}}\rangle}\mathinner{\langle{\mathbf{z}}|}\}$, where $\mathinner{|{\mathbf{z}}\rangle}$ accounts for all possible particle configurations in the probe. For instance, in a probe of size $L$ with $N$ particles (i.e. excitations), there are (${L\choose N}$ different configurations. In the case of $N{=}1$, the problem reduces to the position measurement, discussed above. Each configuration is thus observed by the probability
$p_{{}_{\mathbf{z}}}(t){=}|\braket{\mathbf{z}}{\Psi(t)}|^{2},$ (10)
for which one can obtain the corresponding the CFI. In Fig. 7(b), we compare the long-time behavior of the CFI and the QFI in a half-filled probe of size $L{=}16$ (with $N{=}L/2$ excitations) initialized in the Neel state as a function of $h/J$. Similar to the case of single excitation, the CFI qualitatively follows the QFI, though with lower values. In the inset, we plot $\mathcal{F}_{C}/\mathcal{F}_{Q}$ as a function of $h/J$. Similar to the single excitation case, it shows that a simple configuration measurement is almost optimal in the extended phase and becomes sub-optimal with $\mathcal{F}_{C}/\mathcal{F}_{Q}\simeq 0.5$ as $h/J$ increases. Again the lowest ratio of $\mathcal{F}_{C}/\mathcal{F}_{Q}$ is obtained around the transition point where our simple measurement scheme cannot fully extract the information from the complex critical quantum state of the system.
9 Maximum Likelihood Estimation
The Cramér-Rao inequality only indicates the lower bound of the precision of the probe using an optimal estimation algorithm with optimal measurements. In order to see the real performance of our probe, we specify a measurement basis and use an estimator. Maximum likelihood estimation (MLE) algorithm is recognized as an optimal estimation algorithm for a large measurement samples. For a given measured data set $X$, the estimation for $h$, denoted $h_{es}$, that maximizes the likelihood can be written in the form of
$h_{es}{=}\mathrm{arg}\leavevmode\nobreak\ \underset{\{h\}}{\mathrm{max}}P(X|h)\leavevmode\nobreak\ ,$ (11)
where $P(X|h)$ is the likelihood.
On the other hand, in Bayesian estimation the estimated value is obtained at the point where the posterior probability distribution $P(h|X){=}P(X|h)P(h)/P(X)$ is maximal, where $P(h)$ is the prior probability distribution, and $P(X)$ is the normalization factor. However, when the prior probability distribution is uniform and the sample size is sufficiently large, both estimators converge to identical results [6, 7] . This is because, in the limit of a large sample size, the central limit theorem ensures that the likelihood function approaches a Gaussian distribution. When the sample size is sufficiently large, this Gaussian distribution becomes sharply peaked, concentrating over a narrow interval within the support of the prior. In this paper, we rely on maximum likelihood estimation and thus the choice of prior becomes irrelevant.
In order to obtain our data set, one needs to specify a measurement basis which we take it to be the position measurement whose outcomes appear with probability $p_{{}_{\mathbf{z}}}(t)$ given in Eq. (10). By repeating the measurement for $\mathcal{M}$ times, each outcome $|\mathbf{z}\rangle$ appears for $n_{{}_{\mathbf{z}}}$ times, such that $\sum_{{}_{\mathbf{z}}}n_{{}_{\mathbf{z}}}{=}\mathcal{M}$. Therefore, the likelihood can be written as
$P(X|h)=\begin{pmatrix}\mathcal{M}\\ n_{\mathbf{0}},n_{\mathbf{1}},...,\end{pmatrix}\prod_{\mathbf{z}}p_{\mathbf{z}}^{n_{\mathbf{z}}}\leavevmode\nobreak\ .$ (12)
(a)Refer to caption(b)
Refer to caption(c)Refer to caption(d)
Figure 8: The estimated gradient field $h_{es}/J$ using the maximum likelihood estimator as a function of its true value for sample sizes of $\mathcal{M}=100$ is shown for: (a) the single-particle probe with the system size $L{=}16$; and (b) the multi-particle probe of size $L{=}6$ and $\Delta{=}0$, initialized with the half-filled Neel state. The corresponding standard deviations $\delta h$ are plotted for: (c) single-particle probe, and (d) multi-particle probe. The associated $\delta h$ of the long-time CFI, $1/\sqrt{\mathcal{M}\mathcal{F}_{C}}$, (dashed lines) are depicted for both probes.
We generate random samples based on the true probability distribution of a given $h$. Using the MLE algorithm, $h_{es}$ is taken to be the point at which the likelihood takes its maximum. The uncertainty of the estimation can be quantified by the standard deviation $\delta h$ which is computed over the results of several repetitions of the procedure. In Figs. 8(a)-(b) we show $h_{es}/J$ as a function of real $h/J$ for $\mathcal{M}{=}100$ samples in a single-excitation probe of $L{=}16$ and a multi-particle probe of $L{=}6$ and $N{=}3$ (i.e. half-filling), respectively. As the figures show, the estimation procedure works very well in both single- and multi-particle probes. In Figs. 8(c)-(d), we depict the corresponding standard deviation $\delta h/J$ for the sample size $\mathcal{M}{=}100$ as a function of $h/J$ for the single and the multi-particle probes, respectively. For comparison, we also plot the Cramér-Rao bound $1/\sqrt{\mathcal{M}\mathcal{F}_{C}}$ which shows that the choice of the maximum likelihood estimator can closely converge with the theoretical bound for samples as small as $\mathcal{M}=100$. As shown in Figs. 8(c)-(d), even for $\mathcal{M}{=}100$, the uncertainty closely approaches the theoretical Cramér-Rao bound, indicating near-asymptotic performance and near-saturation of the bound. By further increasing $\mathcal{M}$ the uncertainty decreases and gets closer to the bound.